OpenStudy (baseballer2014):

Rohit is on his way to visit your Grandma, who lives at the end of the state.It's her birthday, and he want to give her the cakes that he has made.Between his place and her grandma house, he need to cross 7 toll bridges. Before you can cross the toll bridge, you need to give them half of the cakes you are carrying, but as they are kind trolls, they each give you back a single cake. How many cakes do rohit have to carry with him so he can reach his grandma home with exactly 2 cakes?

OpenStudy (baseballer2014):

@dan815 @ganeshie8

OpenStudy (dan815):

we can write a function for n tolls

OpenStudy (dan815):

dy/dx=y/2+1

OpenStudy (dan815):

y(7)=2

OpenStudy (dan815):

let x be the number of tolls and y be the number of cakes

OpenStudy (dan815):

we solving for y(0)

OpenStudy (dan815):

not sure if using a continous function is valid though since its a discrete game

OpenStudy (dan815):

another simple way is to go backwards

ganeshie8 (ganeshie8):

\[a_{n+1} = \dfrac{a_n}{2} + 1\]

ganeshie8 (ganeshie8):

it looks like geometric series...

OpenStudy (dan815):

oh 2 cakes!??

OpenStudy (dan815):

u tricky bunny

OpenStudy (dan815):

is the answer just 2 cakes???

OpenStudy (dan815):

if he starts with 2 cakes no matter what hte toll, he can maintain 2 cakes at every end of hte toll

OpenStudy (baseballer2014):

yes

OpenStudy (dan815):

there are other solutions too though probably

OpenStudy (baseballer2014):

i have another problem if you alls re interested more probability

Parth (parthkohli):

lol, dan.

Parth (parthkohli):

Whoops, they return one cake. In that case, we'll have the + sign instead of the - sign.